Determine the equivalent circulating density in ppg for laminar flow of drilling fluid with the following data.. mud weight=9.5ppg, tita 300 =40, tita600=60; plastic viscosity=20cPs; circulation rate=650gpm; hole diameter=8.5 in; drill collect OD=6.75in drill pipe OD=5.0 in; drill collar length=600ft; drill pipe length=10,000ft; true vertical deepth= 9000ft

1
Expert's answer
2021-11-04T00:41:35-0400

In the case of the water stored in a tank, the pressure at its bottom is the weight acting on a unit area of the surface where the tank is kept. To translate that into an equation:

"Pressure = weight/area, and \space weight = mass (m) \times acceleration\space due\space to \space gravity (g)."

This means pressure = m * g/ area.

So we will choose the less circulating density